ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

10. Partial Differential Equations

1991 MS Classification: 46, 47, 81 Chebotarev, Alexander M. Moscow State University, Russia Resolvent theory for quantum dynamical semigroup Quantum dynamical semigroup (QDS) is a solution of the Markov master equation. We describe all conservative (unital) extensions of the minimal QDS introduced by E. B. Davis. The parameterization of extensions is based on resolvent theory for QDS and on Krein-Milman theorem. Reference: A. M. Chebotarev, Math. Notes 56 (1994), N.5-6, 1271–1282.